# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v1

@@ 6.4 Spread option pricing under the fitted model
Finally, over the 1000 simulations, the results from using the fitted model to price spread options are presented in Table 3, with the calibration to forward and call prices shown separately. The spread option prices are computed by the FFT method with $N=2^{8}$, $\bar{u}=40$, which gives sufficient accuracy as noted in Section 6.1. The mean price and its RRMSE are given.
@@
Table 3: Spread option price for the fitted models and their RRMSE over various strike prices $K_{i}$ computed by the FFT method with $N=2^{8}$, $\bar{\theta}=40$. The true prices are from Table 1.
The results show that the spread options are priced less accurately when calibrated to forwards, with an average RRMSE of 30.8%, than to calls, where it is 18.0%. This shows how errors in fitting the model, from 2001 observations and 20 calibration instruments, propagate to pricing errors. As the sample size increases, so that the model parameters are estimated more accurately, the pricing error should approach 0.
It should be unsurprising that calibrating to calls leads to lower pricing error since the forwards calibrate over various maturities, whereas the calls calibrate to the same maturity and the different strikes, which implies out, in part, the marginal component and joint distribution of ${\bf S}(T_{o}){\,|\,}{\cal F}_{T_{e}}$. Both calibration methods produce prices that are downward biased. Surprisingly, calibrating to calls has slightly larger bias and substantially lower standard error here.
@@ Appendix A Choice of model parameters
The parameters in Section 5.1 are chosen to approximately resemble two examples of real data on energy prices appearing in the literature. The first marginal component based on the real data example in (benthbook, , Section 5.1), where $S(t)$ satisfies (4.1) with
$\log S(t)=\Lambda(t)+X(t),$
where $X(t)$ is a LDOUP ${\rm d}X(t)-\lambda X(t){\rm d}t+{\rm d}\widetilde{Z}(t)$. The speed of mean reversion is $\widetilde{\lambda}=0.073$ under their time scale where $t=250$ represents 1 year, and changing to our time scale where $t=1$ does implies $\lambda=\widetilde{\lambda}/\Delta$, where $\Delta=\frac{1}{250}$, with no change to the distribution of $\widetilde{Z}$. This follows from the fact that if ${\bf X}\sim{OU}(\lambda,{\bf Z})$, then $({\bf X}(ct))_{t\geq 0}\sim{OU}(c\lambda,{\bf Z})$ for $c>0$. Thus, we obtain $\lambda$ in (5.5).
The BDLP is a normal inverse Gaussian process $\widetilde{Z}\sim\operatorname{NIG}^{1}(\alpha,\beta,\gamma,\eta)$ with the usual parameterization (see (sco, , Section 5.3.8), $\eta$ here is $\mu$ there). Specifically, their model has $(\alpha,\beta,\gamma)=(4.869,-0.125,0.0874)$, and we set $\eta=0.00284$ to make $\mathbb{E}[\widetilde{Z}(1)]=0$. The VG distribution that matches the 2nd, 3rd, 4th central moments is $\widetilde{Z}\sim VG^{1}(0.42485,\allowbreak-0.00224,0.01796)$. Finally, noting that $\widetilde{Z}=Z(\lambda t)$, we obtain the parameter $\alpha_{1},\mu_{1},\Sigma_{11}$ in (5.5). Similarly, preliminary seasonality parameters $(b_{01},\dots,b_{31})$ can be obtained to exactly match theirs, accounting for the change in time scale.
@@
Energy prices tend to have very high correlation. Accordingly, the joint parameters $a,\Sigma_{12}$ are are chosen so that the Brownian motion correlation is $\rho={\Sigma_{12}}/\sqrt{\Sigma_{11}\Sigma_{22}}=0.95$, and $a=0.9(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}})$. Both are close to the maximum permitted by the parameter constraint $\rho\in(-1,1)$ and $a\in(0,(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}}))$, so that $\operatorname{Corr}(Z_{1}(1),Z_{2}(1))=0.50303$ close to the maximum subject to the marginal components specified above. To obtain a correlation approximately 1 for a WVAG process ${\bf Z}$ would require both $Z_{1}$ and $Z_{2}$ to be approximately equal in law.
@@
where $A$ is the amplitude and $\phi\in[0,1]$ is the phase. The phase is arbitrary and it is reasonable, but not necessary, to expect both energy prices have to have the same phase, so we convert the preliminary seasonal parameters to have $\phi=0.22$ while keeping $A$ unchanged. The intercept $b_{0k}$ affects the price level and may also be considered arbitrary. We set it such that $\boldsymbol{\Lambda}(T_{e})=\log{\bf S}(T_{e})=\log(100,96)$, which is the current energy price. This choice corresponds to assuming ${\bf S}(T_{e})$ is such that ${\bf X}(T_{e})$ is at its stationary mean, with different choices of $b_{0k}$ corresponding to assuming it is above or below that mean. This leads to seasonal parameters in (5.1)–(5.2).
The market price of risk being set to ${\bf h}=(-0.1,-0.03)$ is arbitrary but appears to be within the range of plausible values. See, for example, (BenthPir18, , Section 4.1) in the context of wind power futures.

### v2

@@ 6.4 Spread option pricing under the fitted model
Finally, over the 1000 simulations, the results from using the fitted model to price spread options are presented in Table 3, with the calibration to forward and call prices shown separately. The spread option prices are computed by the FFT method with $N=2^{8}$, $\bar{u}=40$, which gives sufficient accuracy as noted in Section 6.1. The mean prices and their RRMSEs are given.
@@
Table 3: Spread option prices for the fitted models and their RRMSE over various strike prices $K_{i}$ computed by the FFT method with $N=2^{8}$, $\bar{\theta}=40$. The true prices are from Table 1.
The results show that the spread options are priced less accurately when calibrated to forwards, with an average RRMSE of 30.8%, than to calls, where it is 18.0%. This shows how errors in fitting the model, from 2001 observations and 40 calibration instruments, propagate to pricing errors. As the sample size increases, so that the model parameters are estimated more accurately, the pricing error should approach 0.
It should be unsurprising that calibrating to calls leads to lower pricing error since the forwards calibrate over various maturities, whereas the calls calibrate to the same maturity and different strikes, which implies, at least in part, the marginal component distribution and joint distribution of ${\bf S}(T_{o}){\,|\,}{\cal F}_{T_{e}}$. Both calibration methods produce prices that are downward biased. Calibrating to calls has slightly larger bias and substantially lower standard error here.
@@ Appendix A Choice of model parameters
The parameters in Section 5.1 are chosen to approximately resemble two examples of real data on energy prices appearing in the literature. The first marginal component is based on the real data example in (benthbook, , Section 5.1), where $S$ satisfies (4.1) with $\log S(t)=\Lambda(t)+X(t)$, where $X$ is a LDOUP ${\rm d}X(t)-\lambda X(t){\rm d}t+{\rm d}\widetilde{Z}(t)$. The speed of mean reversion is $\widetilde{\lambda}=0.073$ under their time scale where $t=250$ represents 1 year, and changing to our time scale where $t=1$ does implies $\lambda=\widetilde{\lambda}/\Delta$, where $\Delta=\frac{1}{250}$, with no change to the distribution of $\widetilde{Z}$. This follows from the fact that if ${\bf X}\sim{OU}(\lambda,{\bf Z})$, then $({\bf X}(ct))_{t\geq 0}\sim{OU}(c\lambda,{\bf Z})$ for $c>0$. Thus, we obtain $\lambda$ in (5.5).
The BDLP is a normal inverse Gaussian process $\widetilde{Z}\sim\operatorname{NIG}^{1}(\alpha,\beta,\gamma,\eta)$ with the usual parameterization (see (sco, , Section 5.3.8), $\eta$ here is $\mu$ there). Specifically, their model has $(\alpha,\beta,\gamma)=(4.869,-0.125,0.0874)$, and we set $\eta=0.00284$ to make $\mathbb{E}[\widetilde{Z}(1)]=0$. The VG distribution matching the 2nd, 3rd, 4th central moments is $\widetilde{Z}\sim VG^{1}(0.42485,\allowbreak-0.00224,0.01796)$. Finally, noting that $\widetilde{Z}=Z(\lambda t)$, we obtain the parameters $\alpha_{1},\mu_{1},\Sigma_{11}$ in (5.5). Similarly, preliminary seasonality parameters $(b_{01},\dots,b_{31})$ can be obtained to exactly match theirs, accounting for the change in time scale.
@@
Energy prices tend to have very high correlation. Accordingly, the joint parameters $a,\Sigma_{12}$ are chosen so that the Brownian motion correlation is $\rho={\Sigma_{12}}/\sqrt{\Sigma_{11}\Sigma_{22}}=0.95$, and $a=0.9(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}})$. Both are close to the maximum permitted by the parameter constraint $\rho\in(-1,1)$ and $a\in(0,(\frac{1}{\alpha_{1}}\wedge\frac{1}{\alpha_{2}}))$, so that $\operatorname{Corr}(Z_{1}(1),Z_{2}(1))=0.50303$ close to the maximum subject to the marginal components specified above. To obtain a correlation approximately 1 for a WVAG process $\bf Z$} would require both $Z_{1}$ and $Z_{2}$ to be approximately equal in law.
@@
where $A$ is the amplitude and $\phi\in[0,1]$ is the phase. The phase is arbitrary and it is reasonable, but not necessary, to expect both energy prices have the same phase, so we convert the preliminary seasonal parameters to have $\phi=0.22$ while keeping $A$ unchanged. The intercept $(b_{01},b_{02})$ affects the price level and may also be considered arbitrary. We set it such that $\boldsymbol{\Lambda}(T_{e})=\log{\bf S}(T_{e})=\log(100,96)$, which is the current energy price. This choice corresponds to assuming ${\bf S}(T_{e})$ is such that ${\bf X}(T_{e})$ is at its stationary mean, with different choices of $(b_{01},b_{02})$ corresponding to assuming it is above or below that mean. This leads to seasonal parameters in (5.1)–(5.2).
The MPR being set to ${\bf h}=(-0.1,-0.03)$ is arbitrary but appears to be within the range of plausible values. See, for example, (BenthPir18, , Section 4.1) in the context of wind power futures.